A parabola has its vertex at (-3, 2) and its focus at (-3, -1). Write the equations for the parabola, the directrix and the axis of symmetry.
I know about the axis of symmetry and how to get it. I just need to know the other stuff.

Answer by lwsshak3(11628)

A parabola has its vertex at (-3, 2) and its focus at (-3, -1). Write the equations for the parabola, the directrix and the axis of symmetry.
I know about the axis of symmetry and how to get it. I just need to know the other stuff.
**
Use standard form of equation for parabola: (x-h)^2=-4p(y-k), (h,k) being the (x,y) coordinates of the vertex. Parabola opens downwards because focus is below vertex on the axis of symmetry.
For given problem:
vertex: (-3,2) given
axis of symmetry: x=-3
p=distance from vertex to focus on the axis of symmetry=2-(-1)=3
4p=12
Directrix: y=2+p=5
Equation:
(x+3)^2=-12(y-2)